# Thermodynamics of Modified Chaplygin Jacobi And Modified Chaplygin Abel Gas: Stability Analysis and Observational Constraints

### v1

@@ 4   Observational Analysis
In this section of the paper, we outline the observational datasets and the methodology employed to constrain the free parameters of the Modified Chaplygin Jacobi Gas (MCJG) and Modified Chaplygin Abel Gas (MCAG) models. We utilized the Markov Chain Monte Carlo (MCMC) algorithm, specifically employing the emcee package [59]. The likelihood function was constructed assuming Gaussian errors in the observational data, and uniform priors were applied to the model parameters based on physically motivated boundaries.
The MCMC simulations were executed with a total of 10,000 iterations, following an initial burn-in phase of 100 steps to ensure proper convergence of the chains. We employed 100 walkers to efficiently explore the parameter space. The convergence of the MCMC chains was assessed using the Gelman-Rubin statistic, with values of $R<1.1$ for all parameters signifying adequate convergence. The optimal parameter values obtained from the analysis are presented in Table 1, while the posterior distributions at the $1\sigma$ and $2\sigma$ confidence levels are illustrated in Figure 9.
@@
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right]$
@@
$G_{\mu\nu}=8\pi GT_{\mu\nu}$
where the geometric properties of spacetime are encoded in the Einstein tensor $G_{\mu\nu}$, and the matter-energy content is described by the stress-energy tensor $T_{\mu\nu}$.
When we incorporate the FLRW metric into these equations, we obtain the foundational Friedmann equations:
$3H^{2}$ $=\kappa(\rho_{m}+\rho_{de})$ (94)
$3H^{2}+2\dot{H}$ $=-\kappa(p_{m}+p_{de})$ (95)
@@
$\Omega_{\text{total}}=\Omega_{m}+\Omega_{de}=\frac{\rho}{\rho_{\text{crit}}}$
@@
In this context, $\omega_{m}$ represents the constant equation of state (EoS) parameter for the matter component.
Using Eq. (94), the Hubble parameter can be expressed in terms of $\Omega_{m}$ and $\Omega_{de}$ as follows:
$H(z)=H_{0}\left[\Omega_{m0}(1+z)^{3(1+\omega_{m})}+\rho_{de}\right]^{\dfrac{1}{2}}$ (101)
@@
$H(z)_{\text{MCJG}}=H_{0}\left[\Omega_{m0}(1+z)^{3(1+\omega_{m})}+(1-\Omega_{m0})\left(1+2K\left(\dfrac{V_{0}}{V}\right)^{(\alpha+1)(A+1)}\right)^{\frac{1}{1+\alpha}}\right]^{\dfrac{1}{2}}$ (102)
$H(z)_{\text{MCAG}}=H_{0}\left[\Omega_{m0}(1+z)^{3(1+\omega_{m})}+(1-\Omega_{m0})\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\dfrac{1}{2}}$ (103)
@@ 4.1   Cosmic Chronometers (CC) and Baryon Acoustic Oscillations (BAO) Dataset
BAO data provides a standard ruler for measuring the angular diameter distance and the Hubble parameter at different redshifts, based on the imprint of sound waves in the early universe. The BAO dataset employed in this study consists of 26 measurements, including isotropic and anisotropic data from galaxy surveys such as SDSS [66, 67, 68, 69, 70, 71], DES [72], DECaLS [73], and 6dFGS [74]. These measurements cover a redshift range of $0.106\leq z\leq 2.36$. The combined CC+BAO likelihood is computed as:
$\chi^{2}_{\text{CC+BAO}}=-\frac{1}{2}\sum_{i}\left(\frac{H_{\text{obs}}(z_{i})-H_{\text{model}}(z_{i})}{\sigma_{i}}\right)^{2},$
where $H_{\text{obs}}(z_{i})$ represents the observed Hubble parameter at redshift $z_{i}$, $H_{\text{model}}(z_{i})$ denotes the predicted value from the model, and $\sigma_{i}$ indicates the associated uncertainty.
@@ 4.2   Pantheon+SH0ES Supernova Dataset
$E(z^{\prime})=\frac{H(z^{\prime})}{H_{0}}$
@@
Model Parameter Pantheon+SH0ES OHD+BAO
@@ 4.3   Information Criteria
$\text{AIC}=-2\ln(L_{\text{max}})+2k+\frac{2k(k+1)}{N_{\text{tot}}-k-1}$
In this expression, $L_{\text{max}}$ represents the maximum likelihood value, $k$ quantifies the number of model parameters, and $N_{\text{tot}}$ indicates the total quantity of observational datapoints.
The BIC expression incorporates a complexity penalty that scales with sample size:
$\text{BIC}=-2\ln(L_{\text{max}})+k\ln(N_{\text{tot}})$
where the variables maintain their previous definitions: $L_{\text{max}}$ for maximum likelihood, $k$ for parameter count, and $N_{\text{tot}}$ for the observational dataset size.
The DIC computation involves multiple components and is expressed as:
$\text{DIC}=\bar{D}+p_{D}$
@@
$p_{D}=\bar{D}-D(\hat{\theta})$
with $D(\hat{\theta})$ representing the deviance evaluated at the posterior parameter means.
To compare the different models, we compute the AIC, BIC, and DIC for each model and for the $\Lambda$CDM model. The results are presented in Table 2. For comparison purposes, we calculate the relative difference in the Information Criterion (IC) values, defined as:
$\Delta IC_{\text{model}}=IC_{\text{model}}-IC_{\text{min}},$
@@ 4.4   Results and Findings
We evaluate the effectiveness of our models by comparing them against key cosmological data such as the Pantheon+SH0ES dataset. The Hubble parameter values for the MCJG and MCAG models are as follows: for MCJG, $H_{0}=68.170^{+3.585}_{-5.526}$ km s-1 Mpc-1 (Pantheon+SH0ES) and $H_{0}=67.111^{+3.584}_{-5.041}$ km s-1 Mpc-1 (CC+BAO); for MCAG, $H_{0}=72.045^{+0.748}_{-2.659}$ km s-1 Mpc-1 (Pantheon+SH0ES) and $H_{0}=68.842^{+1.572}_{-3.520}$ km s-1 Mpc-1 (CC+BAO). These values are in as per with the estimates obtained from the Planck Collaboration [81], while MCJG exhibit slight deviations from the value obtained by the SH0ES collaboration ($73.24\pm 1.74$ km s-1 Mpc-1) in 2019 [82], MCAG is consistent with it.
@@
The analysis of the Apparent Magnitude vs Redshift plots in Fig. 13 for the Modified Chaplygin Jacobi Gas (MCJG) and the Modified Chaplygin Abel Gas (MCAG) models reveals remarkable similarities in their overall trends, with both showing an increase in apparent magnitude with increasing redshift. This behavior aligns with cosmological expectations, where more distant objects appear fainter. Both models demonstrate strong compatibility with the Pantheon+SHOES observational data, particularly at higher redshifts $(z>0.1)$. However, subtle differences emerge in the low (z < 0.01) and intermediate $(0.01<z<0.1)$ redshift regimes. The MCAG model exhibits a marginally better fit for the nearest objects, while the MCJG model appears to provide a slightly closer fit in the intermediate range. These nuanced distinctions suggest that both models effectively capture the essential features of cosmic acceleration and the transformation from matter-dominated to dark energy-dominated epochs.
@@
The Absolute Magnitude vs Redshift plots for the Modified Chaplygin Jacobi Gas (MCJG) in Fig. 14(a) and the Modified Chaplygin Abel Gas (MCAG) in Fig. 14(b) reveal important insights into the cosmic distance-redshift relationship and the expansion history of the universe. Both plots display the Pantheon+SHOES data as blue points with error bars, showing a wide spread of absolute magnitudes across the redshift range. The best-fit models, represented by green (MCJG) and red (MCAG) lines respectively, demonstrate subtle but significant differences in their predictions. At low redshifts ($z<0.01$), both models show a slight upward trend in absolute magnitude. In the intermediate redshift range ($0.10.01<z<$ 0.1), the models predict a nearly constant absolute magnitude, with the MCJG model showing a slightly flatter trend. At high redshifts (z > 0.1), both models predict a gradual decrease in absolute magnitude, with the MCAG model exhibiting a steeper decline beyond $z\approx 1$.
@@
From Table 1, we observe several key differences and similarities in the parameter values between the MCJG and MCAG models for both the Pantheon+SH0ES and OHD+BAO datasets. For the matter density parameter at the current epoch, $\Omega_{m0}$, the MCJG model estimates it as $0.262_{-0.058}^{+0.023}$ from Pantheon+SH0ES and $0.235_{-0.043}^{+0.034}$ from CC+BAO. In contrast, the MCAG model shows slightly higher values, with $\Omega_{m0}=0.286_{-0.020}^{+0.011}$ (Pantheon+SH0ES) and $0.266_{-0.024}^{+0.024}$ (CC+BAO). These differences suggest that the MCAG model generally assumes a higher matter density at the present epoch compared to the MCJG model, but both are within reasonable ranges compared to standard cosmological models and observational data[83].
The equation of state parameter for matter, $\omega_{m}$, reflects similar behavior in both models. For the MCJG model, $\omega_{m}=0.087_{-0.061}^{+0.067}$ from Pantheon+SH0ES and $0.016_{-0.017}^{+0.035}$ from CC+BAO. The MCAG model exhibits a slightly larger $\omega_{m}$ for Pantheon+SH0ES, $\omega_{m}=0.152_{-0.045}^{+0.055}$, while the CC+BAO value is close to zero at $\omega_{m}=0.011_{-0.028}^{+0.029}$. These results indicate that both models generally assume a nearly dust-like behavior for matter, though the MCAG model suggests a higher equation of state parameter for matter than the MCJG model, especially when using the Pantheon+SH0ES dataset.
The $K$ parameter for the MCJG model is constrained as $K=0.353_{-0.176}^{+0.108}$ (Pantheon+SH0ES) and $K=0.221_{-0.142}^{+0.123}$ (CC+BAO). This parameter does not have an equivalent in the MCAG model, but the slightly higher values from Pantheon+SH0ES suggest that this parameter plays a more prominent role in the MCJG model, particularly in scenarios involving higher redshift data. This value is in close agreement with the previous observational analysis done be Debnath [44].
In the MCAG model, the parameter $e$ is constrained as $e=0.071_{-0.043}^{+0.097}$ from Pantheon+SH0ES and $e=0.088_{-0.059}^{+0.076}$ from CC+BAO. These relatively low values suggest that $e$ has a smaller impact on the overall evolution of the universe in the MCAG model compared to $K$ in the MCJG model. The parameter $c$ in the MCAG model is constrained as $c=0.346_{-0.135}^{+0.125}$ from Pantheon+SH0ES and $c=0.376_{-0.145}^{+0.080}$ from CC+BAO. These values show consistency across datasets, indicating that $c$ plays a comparable role in both Pantheon+SH0ES and CC+BAO observations. Both the constrained values of $e$ and $c$ are in alignment with [44].
Both models include the parameter $\alpha$, though the signs of $\alpha$ differ. For the MCJG model, $\alpha=-0.383_{-0.181}^{+0.252}$ (Pantheon+SH0ES) and $\alpha=-0.451_{-0.174}^{+0.220}$ (CC+BAO), while for the MCAG model, $\alpha=0.576_{-0.372}^{+0.301}$ (Pantheon+SH0ES) and $\alpha=0.483_{-0.332}^{+0.282}$ (CC+BAO). The positive values of $\alpha$ in the MCAG model contrast with the negative values in the MCJG model, suggesting different dynamics in the models regarding the influence of this parameter. The small values of $\alpha$ are in close agreement with the previous works on Chaplygin gas [44, 84]. The negative value of $\alpha$ for the MCJG shows its resemblance with the Generalised Chaplygin gas (GCG) model where a increasingly small and negative value of $\alpha$ is observationally preferred as seen in ref. [85, 84].
For the parameter $A$ the MCJG model gives $A=0.334_{-0.163}^{+0.193}$ (Pantheon+SH0ES) and $A=0.300_{-0.130}^{+0.264}$ (CC+BAO). The MCAG model exhibits higher values, with $A=1.386_{-0.618}^{+0.310}$ (Pantheon+SH0ES) and $A=1.184_{-0.420}^{+0.576}$ (CC+BAO), indicating a more significant influence of the potential energy term in the MCAG model compared to the MCJG model. Though the value of parameter $A$ is in close agreement with [44] for MCJG model but it completely differs in the case of MCAG.
Finally, the parameter $V_{0}$, which is the volume parameter, differs substantially between the two models. For the MCJG model, $V_{0}=0.933_{-0.296}^{+0.495}$ (Pantheon+SH0ES) and $V_{0}=1.064_{-0.369}^{+0.406}$ (CC+BAO). In the MCAG model, however, $V_{0}$ is much larger, with $V_{0}=14.959_{-1.986}^{+2.186}$ (Pantheon+SH0ES) and $V_{0}=15.039_{-1.801}^{+2.009}$ (CC+BAO). This significant difference suggests that the MCAG model assumes a much stronger contribution from the dark energy potential than the MCJG model, which may lead to different cosmic evolution predictions in each scenario.
Lastly, the age of the universe, $t_{\text{age}}$, is found to be $13.78^{+0.14}_{-0.17}$ Gyr for the MCJG model and $13.93^{+0.16}_{-0.16}$ Gyr for the MCAG model. These values show consistency across datasets and align well with the current estimates of the universe’s age from observational data, which support an age of approximately $13.8$ Gyr [83].

### v2

@@ 4   Observational Analysis
In this section of the paper, we outline the observational datasets and the methodology employed to constrain the free parameters of the Modified Chaplygin-Jacobi Gas and Modified Chaplygin-Abel Gas models. We utilized the Markov Chain Monte Carlo (MCMC) algorithm, specifically employing the emcee package [78]. The likelihood function was constructed assuming Gaussian errors in the observational data, and uniform priors were applied to the model parameters based on physically motivated boundaries.
The MCMC simulations were executed with a total of 10,000 iterations, following an initial burn-in phase of 100 steps to ensure proper convergence of the chains. We employed 100 walkers to efficiently explore the parameter space. The convergence of the MCMC chains was assessed using the Gelman-Rubin statistic, with values of $R<1.1$ for all parameters signifying adequate convergence. The optimal parameter values obtained from the analysis are presented in Table 1, while the posterior distributions at the $1\sigma$ and $2\sigma$ confidence levels are illustrated in Figures 9 and 10.
@@
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right]\quad.$
@@
$G_{\mu\nu}=8\pi GT_{\mu\nu}\quad,$
where the geometric properties of spacetime are encoded in the Einstein tensor $G_{\mu\nu}$, and the matter-energy content is described by the stress-energy tensor $T_{\mu\nu}$. When we incorporate the FLRW metric into these equations, we obtain the foundational Friedmann equations:
$3H^{2}$ $=\kappa(\rho_{m}+\rho_{de}),$ (81)
$3H^{2}+2\dot{H}$ $=-\kappa(p_{m}+p_{de}).$ (82)
@@
$\Omega_{\text{total}}=\Omega_{m}+\Omega_{de}=\frac{\rho}{\rho_{\text{crit}}}\quad,$
@@
In this context, $\omega_{m}$ represents the constant equation of state (EoS) parameter for the matter component. Using Eq. (81), the Hubble parameter can be expressed in terms of $\Omega_{m}$ and $\Omega_{de}$ as follows:
$H(z)=H_{0}\left[\Omega_{m0}(1+z)^{3(1+\omega_{m})}+\rho_{de}\right]^{\frac{1}{2}},$ (88)
@@
$H(z)_{\text{MCJG}}=H_{0}\left[\Omega_{m0}(1+z)^{3(1+\omega_{m})}+(1-\Omega_{m0})\left(1+2K\left(\dfrac{V_{0}}{V}\right)^{(\alpha+1)(A+1)}\right)^{\frac{1}{1+\alpha}}\right]^{\dfrac{1}{2}},$ (89)
$H(z)_{\text{MCAG}}=H_{0}\left[\Omega_{m0}(1+z)^{3(1+\omega_{m})}+(1-\Omega_{m0})\left(1+\frac{e^{2}}{c^{2}}\left(\frac{V_{0}}{V}\right)^{e^{2}(A+1)(\alpha+1)}\right)^{\frac{1}{\alpha+1}}\right]^{\dfrac{1}{2}}.$ (90)
@@ 4.1   Cosmic Chronometers (CC) and Baryon Acoustic Oscillations (BAO) Dataset
BAO data provides a standard ruler for measuring the angular diameter distance and the Hubble parameter at different redshifts, based on the imprint of sound waves in the early universe. The BAO dataset employed in this study consists of 26 measurements, including isotropic and anisotropic data from galaxy surveys such as SDSS [85, 86, 87, 88, 89, 90], DES [91], DECaLS [92], and 6dFGS [93]. These measurements cover a redshift range of $0.106\leq z\leq 2.36$. Our analytical methodology addresses the nuanced statistical properties inherent in our cosmological datasets. The dataset comprises two distinct measurement sets: 31 uncorrelated Cosmic Chronometer observations and 26 correlated Baryon Acoustic Oscillation (BAO) measurements, each characterized by unique statistical properties. To capture the intricate statistical landscape, we implement a sophisticated likelihood estimation framework that differentiates between the independent Cosmic Chronometer measurements and the inherently correlated BAO measurements. Our approach utilizes a generalized statistical methodology encoded through the following mathematical formulation:
$\chi^{2}_{\text{composite}}=\sum_{i}\sum_{j}\left[H_{\text{obs}}(z_{i})-H_{\text{model}}(z_{i})\right],\mathbf{V}^{-1}{ij},\left[H{\text{obs}}(z_{j})-H_{\text{model}}(z_{j})\right].$ (91)
In this formulation, $\mathbf{V}^{-1}_{ij}$ represents the inverse of the comprehensive covariance matrix, explicitly modeling the interdependencies present in BAO measurements. For the Cosmic Chronometer subset, the approach simplifies to a diagonal variance representation:
$\chi^{2}_{\text{CC}}=\sum_{i}\left(\frac{H_{\text{obs}}(z_{i})-H_{\text{model}}(z_{i})}{\sigma_{i}}\right)^{2}.$ (92)
This statistical approach ensures rigorous integration of heterogeneous cosmological observations, meticulously accounting for the distinct statistical characteristics of each measurement type. By differentially treating correlated and independent datasets, we enhance the precision and reliability of our cosmological parameter estimations.
@@ 4.2   Pantheon+SH0ES Supernova Dataset
$E(z^{\prime})=\frac{H(z^{\prime})}{H_{0}}.$
@@
In this work, the absolute magnitude $M_{B}$ was determined using the distance modulus $\mu$ available in the Pantheon+SH0ES dataset. The relationship connecting the absolute magnitude $M_{B}$, the distance modulus $\mu$, and the redshift $z$ is given by:
$M_{B}=\mu-5\log_{10}\left(\frac{z(1+z/2)c}{H_{0}}\right)-25,$ (93)
In this expression, $\mu$ represents the distance modulus obtained from the Pantheon+SH0ES dataset, $z$ denotes the redshift, $c=3\times 10^{5}\,\mathrm{km/s}$ is the speed of light, and $H_{0}=70\,\mathrm{km/s/Mpc}$ is the Hubble constant used for the calculation. The apparent magnitude data from the Pantheon+SH0ES dataset was utilized to compute $M_{B}$. The low-redshift approximation for the luminosity distance was applied, assuming $d_{L}(z)\approx\frac{z(1+z/2)c}{H_{0}}$ for small values of $z$. This approximation simplifies the computation and establishes a direct connection between $\mu$ and $M_{B}$. Theoretical predictions for $M_{B}$ were consistently derived using the same conversion formula and subsequently compared with observational data. Notably, this analysis does not directly rely on the low-redshift absolute magnitude provided in the dataset. Instead, the absolute magnitude $M_{B}$ was calculated using the apparent magnitude and the aforementioned formula. This approach ensures consistency in the determination of $M_{B}$ for both observational and theoretical datasets, thereby enabling accurate comparisons and robust analysis.
@@
Model Parameter Pantheon+SH0ES CC+BAO
@@ 4.3   Information Criteria
$\text{AIC}=-2\ln(L_{\text{max}})+2k+\frac{2k(k+1)}{N_{\text{tot}}-k-1}\quad.$
In this expression, $L_{\text{max}}$ represents the maximum likelihood value, $k$ quantifies the number of model parameters, and $N_{\text{tot}}$ indicates the total quantity of observational data-points. The BIC expression incorporates a complexity penalty that scales with sample size:
$\text{BIC}=-2\ln(L_{\text{max}})+k\ln(N_{\text{tot}})\quad.$
where the variables maintain their previous definitions: $L_{\text{max}}$ for maximum likelihood, $k$ for parameter count, and $N_{\text{tot}}$ for the observational dataset size. The DIC computation involves multiple components and is expressed as:
$\text{DIC}=\bar{D}+p_{D}\quad.$
@@
$p_{D}=\bar{D}-D(\hat{\theta})\quad,$
with $D(\hat{\theta})$ representing the deviance evaluated at the posterior parameter means. To compare the different models, we compute the AIC, BIC, and DIC for each model and for the $\Lambda$CDM model. The results are presented in Table 2. For comparison purposes, we calculate the relative difference in the Information Criterion (IC) values, defined as:
$\Delta IC_{\text{model}}=IC_{\text{model}}-IC_{\text{min}}\quad,$
@@ 4.4   Results and Findings
We evaluate the effectiveness of our models by comparing them against key cosmological data such as the Pantheon+SH0ES dataset. The Hubble constant values for the MCJG and MCAG models are as follows: for MCJG, $H_{0}=68.170^{+3.585}_{-5.526}$ km s-1 Mpc-1 (Pantheon+SH0ES) and $H_{0}=67.111^{+3.584}_{-5.041}$ km s-1 Mpc-1 (CC+BAO); for MCAG, $H_{0}=72.045^{+0.748}_{-2.659}$ km s-1 Mpc-1 (Pantheon+SH0ES) and $H_{0}=68.842^{+1.572}_{-3.520}$ km s-1 Mpc-1 (CC+BAO). The analysis of the Hubble constant ($H_{0}$) values for both MCJG and MCAG models reveals interesting insights into the Hubble tension problem. For MCJG, we observe $H_{0}=68.170^{+3.585}_{-5.526}$ km s-1 Mpc-1 (Pantheon+SH0ES), which aligns closely with the Planck collaboration’s CMB-based measurements[100]. This suggests that MCJG maintains consistency with early-universe observations. However, MCAG yields $H_{0}=72.045^{+0.748}_{-2.659}$ km s-1 Mpc-1 (Pantheon+SH0ES), showing better agreement with local measurements from SH0ES collaboration ($73.24\pm 1.74$ km s-1 Mpc-1)[101]. This divergence between the two models can be physically interpreted as follows:
  • MCJG Behavior: The closer alignment with Planck data suggests that MCJG better preserves the standard cosmological evolution from the early universe. This could indicate that the model maintains stronger coupling between matter and radiation components throughout cosmic evolution, leading to a more traditional expansion history.
  • MCAG Behavior: The higher $H_{0}$ value and consistency with SH0ES measurements suggests that MCAG potentially incorporates mechanisms that allow for additional expansion in the late universe, possibly through modified gravitational dynamics or altered cosmic acceleration phases.
Regarding the Hubble tension, these models present distinct physical mechanisms for addressing the discrepancy:
  1. 1.
    MCJG’s Approach: This model favors the early-universe measurements with $H_{0}=68.170^{+3.585}_{-5.526}$ km s-1 Mpc-1, aligning with Planck observations. The larger uncertainties indicate that the model accommodates potential variations in the cosmic expansion rate through:
    • Possible modifications to the standard matter-radiation coupling in the early universe.
    • Flexibility in the dark energy equation of state, allowing for dynamical evolution without violating the cosmic expansion history constraints.
    • Maintenance of consistency with BAO measurements ($H_{0}=67.111^{+3.584}_{-5.041}$ km s-1 Mpc-1), suggesting compatibility with large-scale structure formation.
  2. 2.
    MCAG’s Approach: With $H_{0}=72.045^{+0.748}_{-2.659}$ km s-1 Mpc-1, this model provides a physical resolution to the tension through:
    • Modified gravitational dynamics that could affect late-time cosmic acceleration.
    • Tighter constraints on parameter space, indicated by smaller uncertainties, suggesting a more specific mechanism for reconciling early and late-universe measurements.
    • A natural transition between the CC+BAO value ($68.842^{+1.572}_{-3.520}$ km s-1 Mpc-1) and the local SH0ES measurements, indicating a physically consistent evolution of the expansion rate.
@@
The analysis of the Apparent Magnitude vs Redshift plots in Fig. 14 for the Modified Chaplygin-Jacobi Gas and the Modified Chaplygin-Abel Gas models reveals remarkable similarities in their overall trends, with both showing an increase in apparent magnitude with increasing redshift. This behavior aligns with cosmological expectations, where more distant objects appear fainter. Both models demonstrate strong compatibility with the Pantheon+SH0ES observational data, particularly at higher redshifts $(z>0.1)$. However, subtle differences emerge in the low (z < 0.01) and intermediate $(0.01<z<0.1)$ redshift regimes. The MCAG model exhibits a marginally better fit for the nearest objects, while the MCJG model appears to provide a slightly closer fit in the intermediate range. These nuanced distinctions suggest that both models effectively capture the essential features of cosmic acceleration and the transformation from matter-dominated to dark energy-dominated epochs.
@@
The Absolute Magnitude vs Redshift plots for the Modified Chaplygin-Jacobi Gas in Fig. 15(a) and the Modified Chaplygin-Abel Gas in Fig. 15(b) reveal important insights into the cosmic distance-redshift relationship and the expansion history of the universe. Both plots display the Pantheon+SH0ES data as blue points with error bars, showing a wide spread of absolute magnitudes across the redshift range. The best-fit models, represented by green (MCJG) and red (MCAG) lines respectively, demonstrate subtle but significant differences in their predictions. At low redshifts ($z<0.01$), both models show a slight upward trend in absolute magnitude. In the intermediate redshift range ($0.10.01<z<$ 0.1), the models predict a nearly constant absolute magnitude, with the MCJG model showing a slightly flatter trend. At high redshifts (z > 0.1), both models predict a gradual decrease in absolute magnitude, with the MCAG model exhibiting a steeper decline beyond $z\approx 1$.
@@
From Table 1, we observe several key differences and similarities in the parameter values between the MCJG and MCAG models for both the Pantheon+SH0ES and CC+BAO datasets. For the matter density parameter at the current epoch, $\Omega_{m0}$, the MCJG model estimates it as $0.262_{-0.058}^{+0.023}$ from Pantheon+SH0ES and $0.235_{-0.043}^{+0.034}$ from CC+BAO. In contrast, the MCAG model shows slightly higher values, with $\Omega_{m0}=0.286_{-0.020}^{+0.011}$ (Pantheon+SH0ES) and $0.266_{-0.024}^{+0.024}$ (CC+BAO). These differences suggest that the MCAG model generally assumes a higher matter density at the present epoch compared to the MCJG model, but both are within reasonable ranges compared to standard cosmological models and observational data[102]. The equation of state parameter for matter, $\omega_{m}$, reflects similar behavior in both models. For the MCJG model, $\omega_{m}=0.087_{-0.061}^{+0.067}$ from Pantheon+SH0ES and $0.016_{-0.017}^{+0.035}$ from CC+BAO. The MCAG model exhibits a slightly larger $\omega_{m}$ for Pantheon+SH0ES, $\omega_{m}=0.152_{-0.045}^{+0.055}$, while the CC+BAO value is close to zero at $\omega_{m}=0.011_{-0.028}^{+0.029}$. These results indicate that both models generally assume a nearly dust-like behavior for matter, though the MCAG model suggests a higher equation of state parameter for matter than the MCJG model, especially when using the Pantheon+SH0ES dataset. The $K$ parameter for the MCJG model is constrained as $K=0.353_{-0.176}^{+0.108}$ (Pantheon+SH0ES) and $K=0.221_{-0.142}^{+0.123}$ (CC+BAO). This parameter does not have an equivalent in the MCAG model, but the slightly higher values from Pantheon+SH0ES suggest that this parameter plays a more prominent role in the MCJG model, particularly in scenarios involving higher redshift data. This value is in close agreement with the previous observational analysis done be Debnath [44]. In the MCAG model, the parameter $e$ is constrained as $e=0.071_{-0.043}^{+0.097}$ from Pantheon+SH0ES and $e=0.088_{-0.059}^{+0.076}$ from CC+BAO. These relatively low values suggest that $e$ has a smaller impact on the overall evolution of the universe in the MCAG model compared to $K$ in the MCJG model. The parameter $c$ in the MCAG model is constrained as $c=0.346_{-0.135}^{+0.125}$ from Pantheon+SH0ES and $c=0.376_{-0.145}^{+0.080}$ from CC+BAO. These values show consistency across datasets, indicating that $c$ plays a comparable role in both Pantheon+SH0ES and CC+BAO observations. Both the constrained values of $e$ and $c$ are in alignment with [44]. Both models include the parameter $\alpha$, though the signs of $\alpha$ differ. For the MCJG model, $\alpha=-0.383_{-0.181}^{+0.252}$ (Pantheon+SH0ES) and $\alpha=-0.451_{-0.174}^{+0.220}$ (CC+BAO), while for the MCAG model, $\alpha=0.576_{-0.372}^{+0.301}$ (Pantheon+SH0ES) and $\alpha=0.483_{-0.332}^{+0.282}$ (CC+BAO). The positive values of $\alpha$ in the MCAG model contrast with the negative values in the MCJG model, suggesting different dynamics in the models regarding the influence of this parameter. The small values of $\alpha$ are in close agreement with the previous works on Chaplygin gas [44, 103]. The negative value of $\alpha$ for the MCJG shows its resemblance with the Generalised Chaplygin gas (GCG) model where a increasingly small and negative value of $\alpha$ is observationally preferred as seen in ref. [69, 103]. For the parameter $A$ the MCJG model gives $A=0.334_{-0.163}^{+0.193}$ (Pantheon+SH0ES) and $A=0.300_{-0.130}^{+0.264}$ (CC+BAO). The MCAG model exhibits higher values, with $A=1.386_{-0.618}^{+0.310}$ (Pantheon+SH0ES) and $A=1.184_{-0.420}^{+0.576}$ (CC+BAO), indicating a more significant influence of the potential energy term in the MCAG model compared to the MCJG model. Though the value of parameter $A$ is in close agreement with [44] for MCJG model but it completely differs in the case of MCAG. Finally, the parameter $V_{0}$, which is the volume parameter, differs substantially between the two models. For the MCJG model, $V_{0}=0.933_{-0.296}^{+0.495}$ (Pantheon+SH0ES) and $V_{0}=1.064_{-0.369}^{+0.406}$ (CC+BAO). In the MCAG model, however, $V_{0}$ is much larger, with $V_{0}=14.959_{-1.986}^{+2.186}$ (Pantheon+SH0ES) and $V_{0}=15.039_{-1.801}^{+2.009}$ (CC+BAO). This significant difference suggests that the MCAG model assumes a much stronger contribution from the dark energy potential than the MCJG model, which may lead to different cosmic evolution predictions in each scenario. Lastly, the age of the universe, $t_{\text{age}}$, is found to be $13.78^{+0.14}_{-0.17}$ Gyr for the MCJG model and $13.93^{+0.16}_{-0.16}$ Gyr for the MCAG model. These values show consistency across datasets and align well with the current estimates of the universe’s age from observational data, which support an age of approximately $13.8$ Gyr [102].